OpenStudy (anonymous):

take a square root of a number and then take the square root of that number and so on upto 30

OpenStudy (anonymous):

so u mean for any N take :) \(\sqrt {\sqrt{\sqrt{\sqrt{\sqrt{ \sqrt{\sqrt{\sqrt {\sqrt{\sqrt{\sqrt{\sqrt{ \sqrt{\sqrt{\sqrt {\sqrt{\sqrt{\sqrt{\sqrt{ \sqrt{\sqrt{\sqrt {\sqrt{\sqrt{\sqrt{\sqrt{ \sqrt{\sqrt{\sqrt{\sqrt{N } }}}}}}}} }}}}}}} }}}}}}} }}}}}}}\)

OpenStudy (anonymous):

its the same to \(\Huge N^{\large \frac{1}{2^{30}} } \)

OpenStudy (anonymous):

then what should i do :D